Please help me solve this question:
"The first of two unknown numbers is one less than the second. Four times the first number is four more than twice the second number. Find both the numbers."

Found 2 solutions by lynnlo, ikleyn:
Answer by lynnlo(4176)
Answer by ikleyn(53751)

.
The first of two unknown numbers is one less than the second. Four times the first number is four more
than twice the second number. Find both the numbers.
~~~~~~~~~~~~~~~~~~~~~~~~~

First number x.

Second number (x+1).


The equation is

    4x = 4 + 2(x+1).


Simplify

    4x = 4 + 2x + 2,

    4x - 2x = 4 + 2,

       2x   =   6

        x   =   6/2 = 3.


ANSWER.  First number is 3.  Second number is 4.

Solved.